# Cluster-Norm for Unsupervised Probing of Knowledge

### v1

@@ 4.1.2 Training and Results
These results show that CCS probes trained using our cluster normalization method achieve an average accuracy of $0.77$, while CRC-TPC achieves $0.81$: both relatively high. In contrast, probes following the original CCS approach without clustering tend to perform only slightly better than random guessing. This indicates that our cluster normalization method effectively identifies and eliminates the unwanted contrastive feature from random words (see Section 2.3). Figure 2 visualizes the top principal components of the contrast pair differences when using two random words, clearly illustrating the saliency of this distracting feature under the original setting (left) versus avoiding this problem through cluster normalization (right). For additional details, and replications using other models or harvesting activations at different layers, see Appendix B.1.
@@ 4.2.2 Training & Results
Our results for CCS are shown in Figure 3. We find that a modified prompt including Alice’s opinion causes the majority of our CCS probes to achieve random accuracy against ground truth labels, when normalizing over the entire dataset (Burns-Norm). Clustering before normalizing over each cluster addresses this issue - we see that the average accuracy is closer to that of the control setting, where the opinion of Alice is not inserted. (PCA visualizations analogous to those in Figure 2) are found in Appendix 14.) The reason for cluster normalization achieving higher accuracy is that our method removes the distracting feature of the opinion of Alice, enabling a CCS probe to more accurately determine the direction of the desired knowledge feature. However, for the simpler CRC-TPC method, results differ only slightly for the two approaches, as can be seen in Table 2.
@@
Interestingly, for Mistral-7B, our results differ from those in (Farquhar et al., 2023). Using the default normalization method on the modified dataset, only a few CCS probes are influenced by the explicit opinion of Alice. However, results for other models we have tested (detailed in Appendix B.2) show that the explicit opinion of Alice is indeed often a distraction for the unsupervised probes using only Burns-Normalization, though not as significantly as reported in (Farquhar et al., 2023).
@@ 4.3.2 Training and Results
Logistic Regression 0.81 0.81 0.81
@@
This experiment illustrates when cluster normalization is and is not helpful. Cluster normalization offers a solution to the issue of distracting features, but does not yield a method of unsupervised probing which is robust to prompt changes i.e., differentiation between general knowledge and simulated knowledge. Experimental results using an alternative dataset (TruthfulQA (Evans et al., 2021)), as well as different language models, can be found in Appendix B.3.
@@ 5 Related Work
It has been shown that language models develop internal representations of the world (Li et al., 2022), with individual concepts often encoded as linear directions in activation space (Elhage et al., 2022; Nanda et al., 2023; Burns et al., 2022; Marks & Tegmark, 2023). Language models can also output false information, even if the encoded knowledge in the activations seems to indicate a correct internal representation of the information (Evans et al., 2021; Azaria & Mitchell, 2023; Campbell et al., 2023). We seek to elicit this latent knowledge (Christiano et al., 2022) in an unsupervised manner. In recent years several methods have been proposed (Burns et al., 2022; Belrose et al., 2023, 2024; Zou et al., 2023; Li et al., 2024), although unsupervised methods can be subject to undesirable biases, as shown by Farquhar et al. (2023). They demonstrate that unsupervised probing techniques, such as those developed in Burns et al. (2022), often identify the most salient features in a dataset, as opposed to knowledge only. These features may not always align with the specific knowledge feature of interest, as described in Section 3. We provide theoretical explanations for some of these issues, and propose a method to eliminate them.
The work we cite in the introduction and background sections focuses on finding a general linear representation of knowledge in the latent space of a language model. While we focus on unsupervised approaches, most work concentrates on supervised ones (Christiano et al., 2022; Marks & Tegmark, 2023). This body of work is part of a more general field of research that aims at ensuring truthfulness of language models, by making sure that what they answer is actually what they believe or follows from reasoning e.g., working with quirky language models or using chain-of-thought reasoning (Turpin et al., 2023; Lyu et al., 2023; Radhakrishnan et al., 2023; Mallen & Belrose, 2023).
@@ 7 Limitations
Our study also highlights limitations of current probing techniques that are not addressed by our method. Specifically, as noted by Farquhar et al. (2023), we find that prompting techniques which can induce a language model into simulating a different quality of knowledge by simulating an agent can still affect our unsupervised probe performance. This is a critical limitation, as we specifically want to elicit the knowledge of the model, not that of some simulated entity. Addressing this limitation is another significant challenge for the research community, as it requires an investigation into the question of whether a language model’s knowledge as its capacity to answer a given question under any prompt differs from simulated knowledge, and whether such a difference could be exploited to increase the reliability of probing algorithms. These limitations are studied in Mallen & Belrose (2023), where the context-dependence of knowledge probes is measured.
@@ A.2 Results
Appendix B Additional Results
In addition to Mistral-7B, the random word experiment and explicit opinion experiment is repeated for the following models: Gemma-7b, Phi-2 and 3, Llama 3 8B and Pythia-6.9b-v0. We harvested activations at four different points: the 25th percentile layer, 50th, 75th and the last layer. Unbiased examples correspond to probes trained on the original prompts, while biased examples correspond to probes trained on the modified prompts.
@@ B.1 Random Words
The results for the additional models and layers are comparable to those of Mistral-7b at the 75th percentile layer. Figure 7 shows the average results across all models, including Gemma-7b, Phi-2 and 3, Llama 3 8B, Pythia-6.9b-v0, and Mistral-7b. For the biased prompt-template dataset, unsupervised methods using Cluster-Normalization do usually perform better than those using the standard Burns-Normalization. Results can be found in figures 8, 9, 10, 11, 12 and 13.
@@ B.2 Explicit Opinion
Figure 14 shows a PCA visualizations of contrast differences without our cluster normalization, as figure 2.
The results for the additional models and layers are comparable to those of Mistral-7b at the 75th percentile layer. Figure 15 shows the average results across all models, including Gemma-7b, Phi-2 and 3, Llama 3 8B, Pythia-6.9b-v0, and Mistral-7b. For the unbiased and especially the biased prompt-template dataset, unsupervised methods using Cluster-Normalization tend to outperform those that use standard Burns-Normalization. However, the performance gap between these two methods is smaller compared to the random word experiment. Moreover, in our work, the standard CCS using Burns-Normalization appears to perform better on the biased dataset than reported by Farquhar et al. (2023) for the different models and layers. The individual results for the various layers and models are shown in the figures: 17, 18, 19, 20 and 21.
Figure 7: Mean accuracy of Logistic Regression, CRC and CCS probes averaged across all models on original templates (up) and biased ones (down) for the random word experiment.
Figure 8: Mean accuracy of Logistic Regression, CRC and CCS probes on Llama-3-8B on original templates (up) and biased ones (down) for the random word experiment.
Figure 9: Mean accuracy of Logistic Regression, CRC and CCS probes on Mistral-7B on original templates (up) and biased ones (down) for the random word experiment.
Figure 10: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-2 on original templates (up) and biased ones (down) for the random word experiment.
Figure 11: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-3 on original templates (up) and biased ones (down) for the random word experiment.
Figure 12: Mean accuracy of Logistic Regression, CRC and CCS probes on Gemma-7b original templates (up) and biased ones (down) for the random word experiment.
Figure 13: Mean accuracy of Logistic Regression, CRC and CCS probes on Pythia-6.9b-v0 original templates (up) and biased ones (down) for the random word experiment.
@@
Figure 15: Mean accuracy of Logistic Regression, CRC and CCS probes on LLama-3-8b on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 16: Mean accuracy of Logistic Regression, CRC and CCS probes on LLama-3-8b on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 17: Mean accuracy of Logistic Regression, CRC and CCS probes on Mistral-7b on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 18: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-2 on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 19: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-3-Instruct Mini on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 20: Mean accuracy of Logistic Regression, CRC and CCS probes on Gemma-7b on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 21: Mean accuracy of Logistic Regression, CRC and CCS probes on Pythia-6.9b-v0 on original templates (up) and biased ones (down) for the explicit opinion experiment.
Figure 22: Mean accuracy of Logistic Regression, CRC and CCS probes averaged across all models on original templates (up) and biased ones (down) for the explicit opinion experiment.
@@ B.2.1 Violin Plots — 75th percentile layer
Figure 23: Llama-3-8b
Figure 24: Phi-3 Mini
@@
Figure 26: Gemma 7b
Figure 27: Pythia 6.9b

### v2

@@ 4.1.2 Training and Results
These results show that CCS probes trained using our cluster normalization method achieve an average accuracy of $0.77$, while CRC-TPC achieves $0.81$: both relatively high. In contrast, probes following the original CCS approach without clustering tend to perform only slightly better than random guessing. This indicates that our cluster normalization method effectively identifies and eliminates the unwanted contrastive feature from random words (see Section 2.3). Figure 3 visualizes the top principal components of the contrast pair differences when using two random words, clearly illustrating the saliency of this distracting feature under the original setting (left) versus avoiding this problem through cluster normalization (right). Figure 2 shows the mean accuracy for different layers across multiple models, including Mistral-7B and additional models Gemma-7B, Phi-2, Phi-3, Llama-3-8B, and Pythia-6.9B-v0. The results demonstrate that our Cluster-Norm method significantly enhances the performance of unsupervised methods.
@@ 4.2.2 Training & Results
Our results for CCS are shown in Figure 4. We find that a modified prompt including Alice’s opinion causes the majority of our CCS probes to achieve random accuracy against ground truth labels, when normalizing over the entire dataset (Burns-Norm). Clustering before normalizing over each cluster addresses this issue - we see that the average accuracy is closer to that of the control setting, where the opinion of Alice is not inserted. (PCA visualizations analogous to those in Figure 3 are found in Appendix 14, while a figure displaying the mean accuracy for different layers across multiple model can be found in Appendix B.2.) The reason for cluster normalization achieving higher accuracy is that our method removes the distracting feature of the opinion of Alice, enabling a CCS probe to more accurately determine the direction of the desired knowledge feature. However, for the simpler CRC-TPC method, results differ only slightly for the two approaches, as can be seen in Table 2.
@@
Interestingly, for Mistral-7B, our results differ from those in Farquhar et al. (2023). Using the default normalization method on the modified dataset, only a few CCS probes are influenced by the explicit opinion of Alice. However, results for other models we have tested (detailed in Appendix B.2) show that the explicit opinion of Alice is indeed often a distraction for the unsupervised probes using only Burns-Normalization, though not as significantly as reported in Farquhar et al. (2023).
@@ 4.3.2 Training and Results
Logistic Regression (Upper Bound) 0.81 0.81 0.81
@@
This experiment illustrates when cluster normalization is and is not helpful. Cluster normalization offers a solution to the issue of distracting features, but does not yield a method of unsupervised probing which is robust to prompt changes i.e., differentiation between general knowledge and simulated knowledge. Experimental results using an alternative dataset (TruthfulQA (Evans et al., 2021)) and different models, can be found in Appendix B.3.
@@ 5 Related Work
It has been shown that language models develop internal representations of the world (Li et al., 2022), with individual concepts often encoded as linear directions in activation space (Elhage et al., 2022; Nanda et al., 2023; Burns et al., 2022; Marks and Tegmark, 2023). Language models can also output false information, even if the encoded knowledge in the activations seems to indicate a correct internal representation of the information (Evans et al., 2021; Azaria and Mitchell, 2023; Campbell et al., 2023). We seek to elicit this latent knowledge (Christiano et al., 2022) in an unsupervised manner. In recent years several methods have been proposed (Burns et al., 2022; Belrose et al., 2023, 2024; Zou et al., 2023; Li et al., 2024), although unsupervised methods can be subject to undesirable biases, as shown by Farquhar et al. (2023). They demonstrate that unsupervised probing techniques, such as those developed in Burns et al. (2022), often identify the most salient features in a dataset, as opposed to knowledge only. These features may not always align with the specific knowledge feature of interest, as described in Section 3. We provide theoretical explanations for some of these issues, and propose a method to eliminate them.
The work we cite in the introduction and background sections focuses on finding a general linear representation of knowledge in the latent space of a language model. While we focus on unsupervised approaches, most work concentrates on supervised ones (Christiano et al., 2022; Marks and Tegmark, 2023). This body of work is part of a more general field of research that aims at ensuring truthfulness of language models, by making sure that what they answer is actually what they believe or follows from reasoning e.g., working with quirky language models or using chain-of-thought reasoning (Turpin et al., 2023; Lyu et al., 2023; Radhakrishnan et al., 2023; Mallen and Belrose, 2023).
@@ 7 Limitations
Our study also highlights limitations of current probing techniques that are not addressed by our method. Specifically, as noted by Farquhar et al. (2023), we find that prompting techniques which can induce a language model into simulating a different quality of knowledge by simulating an agent can still affect our unsupervised probe performance. This is a critical limitation, as we specifically want to elicit the knowledge of the model, not that of some simulated entity. Addressing this limitation is another significant challenge for the research community, as it requires an investigation into the question of whether a language model’s knowledge as its capacity to answer a given question under any prompt differs from simulated knowledge, and whether such a difference could be exploited to increase the reliability of probing algorithms. These limitations are studied in Mallen and Belrose (2023), where the context-dependence of knowledge probes is measured.
@@ A.2 Results
Appendix B Additional Probing Results
In addition to Mistral-7B, the random word experiment and explicit opinion experiment is repeated for the following models: Gemma-7B, Phi-2 and 3, Llama-3-8B and Pythia-6.9B-v0. We harvested activations at four different points: the 25th percentile layer, 50th, 75th and the last layer. Unbiased examples correspond to probes trained on the original prompts, while biased examples correspond to probes trained on the modified prompts.
@@ B.1 Random Words
The results for the additional models and layers are comparable to those of Mistral-7B at the 75th percentile layer. For the biased prompt-template dataset, unsupervised methods using cluster normalization do usually perform better than those using the standard Burns-Normalization. Figures 8, 9, 10, 11, 12, and 13 show violin plots of the results for the additional models.
@@ B.2 Explicit Opinion
Figure 14 shows PCA visualizations of contrast differences without our cluster normalization, analogous to Figure 3.
The results for the additional models and layers are comparable to those of Mistral-7B at the 75th percentile layer. Figure 15 shows the average results across all models, including Gemma-7B, Phi-2, Phi-3, Llama-3-8B, Pythia-6.9B-v0, and Mistral-7B. For both the unbiased and biased prompt-template datasets, unsupervised methods using cluster normalization tend to outperform those using standard Burns-Normalization, with the difference being more pronounced for the biased dataset. However, the performance gap between these two methods is smaller compared to the random word experiment. Moreover, in our work, the standard CCS using Burns-Normalization appears to perform better on the biased dataset than reported by Farquhar et al. (2023) for the various models and layers. The individual results for the different layers and models are shown in the following figures: 17, 18, 19, 20 and 21.
Figure 8: Mean accuracy of Logistic Regression, CRC and CCS probes on Llama-3-8B on original prompts (up) and biased ones (down) for the random word experiment.
Figure 9: Mean accuracy of Logistic Regression, CRC and CCS probes on Mistral-7B on original prompts (up) and biased ones (down) for the random word experiment.
Figure 10: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-2 on original prompts (up) and biased ones (down) for the random word experiment.
Figure 11: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-3 on original prompts (up) and biased ones (down) for the random word experiment.
Figure 12: Mean accuracy of Logistic Regression, CRC and CCS probes on Gemma-7B original prompts (up) and biased ones (down) for the random word experiment.
Figure 13: Mean accuracy of Logistic Regression, CRC and CCS probes on Pythia-6.9B-v0 original prompts (up) and biased ones (down) for the random word experiment.
@@
Figure 15: Mean accuracy of Logistic Regression, CRC-TPC, and CCS probes across six models for (top) original and (bottom) biased prompts in the explicit opinion experiment. CCS probes using Cluster-Normalization consistently outperform those using Burns-Normalization, particularly for modified prompts, across the 25th and 75th percentile layers and the final layer.
Figure 16: Mean accuracy of Logistic Regression, CRC and CCS probes on LLama-3-8b on original prompts (up) and biased ones (down) for the explicit opinion experiment.
Figure 17: Mean accuracy of Logistic Regression, CRC and CCS probes on Mistral-7B on original prompts (up) and biased ones (down) for the explicit opinion experiment.
Figure 18: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-2 on original prompts (up) and biased ones (down) for the explicit opinion experiment.
Figure 19: Mean accuracy of Logistic Regression, CRC and CCS probes on Phi-3-Instruct Mini on original prompts (up) and biased ones (down) for the explicit opinion experiment.
Figure 20: Mean accuracy of Logistic Regression, CRC and CCS probes on Gemma-7B on original prompts (up) and biased ones (down) for the explicit opinion experiment.
Figure 21: Mean accuracy of Logistic Regression, CRC and CCS probes on Pythia-6.9B-v0 on original prompts (up) and biased ones (down) for the explicit opinion experiment.
Figure 22: Mean accuracy of Logistic Regression, CRC and CCS probes averaged across all models on original prompts (up) and biased ones (down) for the explicit opinion experiment.
@@ B.2.1 Violin Plots — 75th Percentile Layer
Figure 23: Llama-3-8B - Explicit Opinion Experiment
Figure 24: Phi-3 Mini - Explicit Opinion Experiment
@@
Figure 26: Gemma-7B - Explicit Opinion Experiment
Figure 27: Pythia-6.9B - Explicit Opinion Experiment
@@ B.3 Prompt Template Sensitivity
Appendix C Zero-Shot Results
For reference, we present the zero-shot results for different models.
C.1 Random Words
For the random word experiment, we report only the zero-shot results using the default prompts, where no random word is appended. This decision stems from the structure of our biased prompts and the nature of zero-shot prediction. Recall that for biased prompts we append a random word after the sentiment positve or negative pseudo-label:
Consider the following example: [review],
Between positive and negative, the
sentiment of this example
is [label]. [random_word]
However, in the zero-shot scenarios, the model would need to predict the label itself. Consequently, using modified prompts for zero-shot prediction becomes impractical here. We therefore limit our reporting to results from the default prompts for this experiment. The results are presented in Table 5. On average, The zero-shot accuracies observed in this experiment are notably lower than the probe acccuracies for the 75th percentile layer, as illustrated in Figure 2.
Model Default
Gemma-7B 0.51
Llama-3-8B 0.76
Mistral-7B 0.86
Phi-2 0.84
Phi-3 0.94
Pythia-6.9B 0.51
Table 5: Random Word experiment: Zero-shot prompting performance for different models on the IMDb dataset for the default prompt templates only (no random word is appended to a prompt).
C.2 Explicit Opinion
We evaluate zero-shot performance using unbiased (standard) and biased (with added explicit opinion) prompt templates across six language models: Gemma-7B, Llama-3-8B, Mistral-7B, Phi-2, Phi-3, and Pythia-6.9B. Table 6 presents the results. Similar to the random words experiment, we observe that the zero-shot accuracies are, on average, lower than the probe acccuracies for both default and modified prompt templates, as illustrated in Figure 22. However, the difference between the zero-shot accuracies for default and modified prompts is not substantial for most models.
Model Default With Explicit Opinion
Gemma-7B 0.51 0.51
Llama-3-8B 0.59 0.60
Mistral-7B 0.68 0.59
Phi-2 0.54 0.68
Phi-3 0.93 0.92
Pythia-6.9B 0.50 0.50
Table 6: Explicit Opinion experiment: Zero-shot performance comparison on the IMDb dataset. Results show accuracies for default prompt templates (no explicit opinion is added) and modified prompt templates (with added explicit opinion).
C.3 Prompt Template Sensitivity
Model Prompt Template
Default Literal Professor
Mistral-7B 0.72 0.71 0.65
Llama-3-8B 0.69 0.64 0.72
Phi-2 0.54 0.59 0.54
Table 7: Zero-shot performance for the Prompt Template Sensitivity experiment.
In this experiment, we evaluate zero-shot performance using the same three templates: Default, Literal, and Professor, tested with Mistral-7B, Llama-3-8B, and Phi-2. The corresponding zero-shot accuracies are shown in Table 7. Notably, we do not necessarily see higher performance using the “Professor” template for all models. While this outcome is unexpected, it is difficult to speculate on its underlying cause; we instead focus on the effect of cluster normalization in our main discussion of results.
C.4 Implicit Opinion
Model Company Non-Company
Unbiased Biased Unbiased Biased
Mistral-7B 0.96 0.39 0.98 0.62
Llama-3-8B 0.97 0.17 0.98 0.53
Phi-2 0.98 0.39 0.94 0.90
Table 8: Zero-shot performance performance for the Implicit Opinion Experiment. Here, the effect of Alice’s biased implicit opinion is clearly demonstrated.
Following the setup of this experiment, we examine zero-shot performance for both “company” and “non-company” questions, tested with Mistral-7B, Llama-3-8B, and Phi-2. Our results are summarized in Table 8.
Focusing on Mistral-7B, but with the same pattern of results apparent for all three models, introducing Alice’s biased implicit opinion substantially decreases zero-shot accuracy for questions labeled as “company”, from 0.96 to 0.39. Conversely, for non-company labels, accuracy declines moderately from 0.98 to 0.62 when the biased setting is applied. These results underscore the significant influence of biased prompts on model performance, particularly in scenarios involving implicit opinions.